A fish takes the bait and pulls on the line with a force of 2.2 N . The fishing reel, which rotates without friction, is a uniform cylinder of radius 0.050 m and mass 0.97 kg . Part A What is the angular acceleration of the fishing reel? Express your answer using two significant figures. Part B How much line does the fish pull from the reel in 0.20 s ? Express your answer using two significant figures.
Answer #1

Moment of inertia of the reel

I = mR^2/2

and
torque

M = RF

angular acceleration
α = Μ/Ι = 2F/mR = 4.4/(0.97*0.050) =90.72 rad/s^2

linear acceleration
a = Rα = 4.53 m/s^2


L = (1/2)at^2 = 0.5*4.53*0.04 = 0.09 m